Two sides of a triangle are 12 cm and 14 cm. The perimeter of the triangle is 36 cm. What is its third side?

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

Let x cm be the third side Perimeter of triangle = 36 cm
12 + 14 + x = 36
26 + x = 36
x = 36 - 26
x = 10 cm
∴The third side is 10 cm

​Thanks
Harshit Singh
3 years ago
Dear Student

Let x cm be the third side Perimeter of triangle = 36 cm
12 + 14 + x = 36
26 + x = 36
x = 36 - 26
x = 10 cm
∴The third side is 10 cm

​Thanks